During March, the production department of a process operations system completed and transferred to finished goods 24,000 units that were in process at the beginning of March and 110,000 units that were started and completed in March. March's beginning inventory units were 100% complete with respect to materials and 45% complete with respect to labor. At the end of March, 27,000 additional units were in process in the production department and were 100% complete with respect to materials and 65% complete with respect to labor. The production department incurred direct materials cost of $254,600 and its beginning inventory included materials cost of $94,000. Compute the direct materials cost per equivalent unit for the department using the weighted-average method.

a. 3.17

b. 2.60

c. 2.17

d. 1.37

e. 2.31

Please show me all the steps if possible.
Answer #1

Equivalent unit of material = 24000+110000+27000 = 161000

Total material cost = 254600+94000 = 348600

Cost per equivalent unit of direct material = 348600/161000 = 2.17

So answer is c) $2.17